Previous

Calendar

June

2023

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

28

JD148
In Lab/Cristol (SpB): Details , Photos (4)

started using AP200_spb05.CR1X (added \r \n to serial output from AP200, this is what I needed!)

booted titan2 in Cristol 222

at 14:45 MST, modified niwot.xml, start of ops105 (added some AP200 variables to titan2)

29

JD149
CU Holiday (Memorial Day)

30

JD150

31

JD151

1

JD152
Site Visit (SpB): Details , Photos (29)

Span cylinder at 1600 psi, N2 at 900 psi

at 9:00 MST, downloaded PolaRx5 GPS "B" (sn 3227662) data files (in forest)

at 9:10 MST, swapped SD card from GPSCAM

at 9:15 MST, reconnected the grounding wire to the forest PolaRx5 GPS "B" reciever

at 10:00 MST, downloaded PolaRx5 GPS "A" (sn 3062189) data files

at around 10:10 MST, hit the "reset" button on the 21.5m LI-7500A (didn't seem to do anything)

drove the road from MRS to C-1 for the first time this year (moved 5 empty cylinders from C-1 to the snowshed)

transported AP200 boxes #1 and #2 from CU campus to our trailer at C-1

Enterprise SUV rental

Overcast, calm, started to rain/hail

2

JD153

3

JD154
(Full "Strawberry" Moon)

Wet

4

JD155
at approx 12:55 MST, the fast cr23x logger data (ch 1,30) stops being ingested by the DSM!

5

JD156
In Lab/Cristol (SpB): Details , Photos (7)

tried to get urquell2 working again (had the wrong iso disk)

at US-NR1: at approx 8:15 MST, the fast cr23x logger data (ch 1,30) started being ingested by the DSM again.

Based on T.soil, snow is now gone at the mini-tower (start of the warm season)

6

JD157
In Lab/Cristol (SpB): Details , Photos (9)

tried to use AP200_spb06.CR1X and turn on the pump (did not seem to work)

able to get urquell2 working again!

7

JD158
In Lab/Cristol (SpB): Details

started using AP200_spb06.CR1X (output more ap200 variables)

at 19:04 MST, modified niwot.xml, start of ops106 (added more AP200 variables as output)

Heavy rain (according to Kris)

8

JD159
Site Visit (SpB, PB): Details , Photos (35)

Span cylinder at 1550 --> 1850 psi, N2 at 800 psi

at 14:00 MST, swapped span cylinder; old=cc477126 (at 1550 psi) new=ff66921 (at 1850 psi)

at 11:00 MST, downloaded PolaRx5 GPS "A" (sn 3062189) data files

pulled in the 21.5m LI-7500A and it started working again (left there for now)

at 12:40 MST, removed the 2.5m LI-7500 (sn 75H-0062) and control box (bring back to campus)

added boom/pulley system at the top of the tower (for the LBL instrument IC)

moved the LBL cal gas cylinders (ff66921 and ff66936) from campus to the site

from 13:19-13:21 MST, adjusted the pressure/flow in the span/n2 tanks (for LI-6262)

Enterprise SUV rental

Overcast, Big clouds around

9

JD160
In Lab/Cristol (SpB): Details , Photos (3)

moved the boxes for LBL comparison from the loading dock to Cristol 222

10

JD161

11

JD162
In Lab/Cristol (SpB, SC): Details

picked up 3 boxes, 3 cylinders for SC/HC

put the 2.5m LI-7500 (sn 75H-0062) and control box in lab

12

JD163
Site Visit (SpB, PB, SC, HC): Details , Photos (32)

Span cylinder at 1650 psi, N2 at 700 psi

moved LBL Tower Team boxes to the tower

started to do setup, but lightning/rain/hail stopped the work

Start of the US-NR1 instrument comparison by the AmeriFlux TowerTeam (SC/HC)

Hail, some rain

13

JD164
Site Visit (SC, HC): Details

SC/HC continue LBL Tower Team setup

Wet?

14

JD165
Flag Day

15

JD166
In Lab/Cristol (SpB): Details , Photos (6)

Now using the RS232 port on the CR1000x to output data to the titan

16

JD167

17

JD168

18

JD169
Fathers Day

19

JD170
Start of the NSF Fluxcourse at MRS.

20

JD171
In Lab/Cristol (SpB): Details

connected the AC power supply to the AP200 (seems to keep working fine)

AP200 box with AC power: photo

made usb backups (see usb_disk_western_digital_my_passport_5Tb.txt)

21

JD172
Site Visit (SpB): Details , Photos (38)

Span cylinder at 1350 psi, N2 at 600 psi

at 10:25 MST, downloaded PolaRx5 GPS "A" (sn 3062189) data files

power for GPS "A"; now using a AC/DC adapater (previously powered from the CNR cr23x logger)

at 11:10 MST, downloaded PolaRx5 GPS "B" (sn 3227662) data files (in forest)

releveled the GPS "B" antenna (see photos)

at 11:15 MST, swapped SD card from GPSCAM

downloaded logger data from the A. Richardson CR800 and CR1000

SpB Prius

Sunny w/ big clouds down on plains

Summer Solstice

22

JD173
Site Visit (SpB, PB, KN, +): Details , Photos (11)

Tour of the US-NR1 site by Fluxcourse.

Sunny in AM, rain/clouds in PM

23

JD174
Site Visit (SpB, SC, HC, JS, KB, A): Details , Photos (19)

End of the US-NR1 instrument comparison by the AmeriFlux TowerTeam (SC/HC)

moved the ap200 out to the tower (in the green shed)

removed all the TowerTeam comparison instruments

Sunny

24

JD175

25

JD176
Site Visit (SpB, JvB, BsB, DD): Details , Photos (22)

Turned on the fan in the LI-6262 box

Nice weather.

26

JD177

27

JD178
Jane S. team setup dust samplers on the tower.

28

JD179

29

JD180
Jane S. team removed dust samplers from the tower.

30

JD181
Site Visit (SpB, WW, DL, +): Details , Photos (23)

Tour of the US-NR1 site and ecoTram by NCAR CGD.

End of the NSF Fluxcourse at MRS.

SpB Prius

Rain held off till 1pm'ish

1

JD182